PartnerObserving Log for 2007 Mar 11/12

Observer: John Hill
TO: Ed Lundin

Some more bonus SDT observations in engineering time. As with last night, I'm only using SDT OBs to test various items. In this case, the photometric conditions may yield some useful data.

10 Biases

Opened at sunset. Crystal clear = photometric. Light wind. T=2degC.

01:35-0148 Short Exposure Tests (sub-rastered images can be collected at intervals of 18-24 sec.)

SX Mirror panic'd -- not sure why.

01:49-02:00 SDT_Uspec sky flats failed because of LBC-Blue shutter communication problem

02:02 Turn off/Turn on LBC to cure shutter problem (lost 20 min).

02:00 Notice control room getting warm, and find that main chillers in Utility Bldg have shut down. Turn off mirror ventilation, HBS, and cell compressors. Then made a trip to utility bldg. Power cycle both chillers and restart. Seems OK as we wait for the loops to cool down. Turn off/on the chiller on level 4 and wait for the cold loop to cool down. Back on the sky about 04:20. See Issue #533 (lost 2:20 hours). P.S. Ed was looking kind of bored last night, so we decided to break something big tonight to keep him enriched.

04:12 Another short exposure test with shutter working -- same result.

04:23-04:30 Bfastextra for WF

04:35 LeoII_1_U nearly all of the dither positions failed to guide for lack of a guide star (All these U exposures are open loop tracking --see below.)

05:05-05:09 Bfastextra for WF after reactivating SX actuators 209 and 328 (they failed to turn on when we restarted the mirror).

05:12 LeoII_2_U -- much clearer than last night. Now we can see the stars in the LeoII dwarf.

05:43 Bfastextra for WF

05:45 LeoII_3_U

06:16 Bfastextra for WF

06:19 LeoII_4_U -- seeing about 1.5 arcsec in U

06:51 Bfastextra for WF

06:56 LeoII_R -- EL=80deg on meridian and longer wavelength bring seeing down to 1.1 arcsec. Guider seems to be struggling in variable seeing, but then we notice that guiding is disabled in PCS. Guiding enabled for dither 5. The good news is that the telescope open-loop trails are less than 1.5 arcsec in 5 minutes at elevation 80deg.

07:27 Bfastextra for WF

07:29 Sfinefoc to check focus for SDT_Uspec filter. Likes offset 0.44 mm in 5.7 pix FWHM, but very few stars in this field.

07:34 repeat LeoII_R -- Guiding much better, but seeing worse. Wind increasing from the NE. Seeing improves toward the end.

08:05 Bfastextra for WF

08:08 Sfinefoc to check focus for SDT_Uspec filter. Likes offset 0.45 mm in 4.5 pix FWHM.

08:13 LeoII_V -- Seeing variable. Guiding failed on dither 4.

I've updated the FITS headers to display the proper units ("nm" rather than "microns") for the Zernike terms from PSF. Note that for now these Zernike terms are the last increment which was applied to the mirror by active optics and NOT the total correction.

08:42 Bfastextra at position of sdt_fan1_3

08:45 sdt_fan1_3 -- Guiding failed on dither 1. SX Mirror panic'd during dither 4. Timeout on getdata from hardpoint 5. Rebooting VXWorks didn't help -- cell program can't talk to hardpoint 5. Cycled power supplies and power to VxWorks in order to restore the cell. See Issue #532 (lost 45 min). Note interesting super-extra-focal image 090846 from dither 4 with the mirror down. P.S. Ed is feeling quite enriched now.

09:37 Resume dither 5 of sdt_fan1_3, now with plenty of 6-day moonlight. The OB positioned the telescope for dither 5, but then it ended the OB without taking an image. (lost 10 min)

09:50 Bfastextra

09:54 sdt_fan1_4

10:24-10:30 Bfastextra at position of A1668 (No good pupils; had to double exposure time; this A1668 field could use dedicated WF OBs with brighter stars)

10:34 A1668_B -- All the exposures were accidentally 2x longer than requested because I forgot to put the scaling back after Bfastextra. Dither 1 is guiding on a disk galaxy with a bright nucleus (What do you expect in an Abell cluster?). B images are nice and round in 1 arcsec seeing - No sign of field dependent astigmatism in this image at EL=64deg. Second image is a bit better after guiding on a star. Dither 3 suffered some kind of a jump or guding error making the images about 2 arcsec long. Guiding failed on dither 4 -- the UI is displaying a box with no hint of a star.

10:54 Bfastextra 32s -- got a bad fit to one of two pupils, so crazy wavefront numbers which I didn't send. Possibly this was because of a bad column going through the pupil.

10:58 A1668_U1 -- Dither 1 guiding on the galaxy again. Dither 2 has a nice sharp star to guide on, but the guiding corrections are very jumpy (at the half arcsec level) - not sure why. Guiding was OK on dither 3, and guiding failed again on dither 4.

11:25 Bfastextra at position of A1668_U2 -- UI froze during this exposure, and then took a second image?? Did it accept my second click of the Play button even though it appeared frozen?

11:30 A1668_U2 -- Dither 1 guiding on that same pesky disk galaxy even though the pointing has changed. Dithers 2,3 had successful guiding. Guiding failed on dither 4.

11:57 Bfastextra -- No good pupils found in 16 sec.

12:04 Bfastextra at 16:00:00 +20:00:00 -- No good pupils found in 16 sec. I take this as a sign that we should go to bed.

12:12 Close Dome. T=-0.3degC Wind=4m/sec from NE D=-9degC

-- JohnHill - 12 Mar 2007
Topic revision: r4 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback